OpenStudy (anonymous):

Solve the system using elimination. 2x+6y=-12 5x-5y=10 A.(2, 1) B.(0, –2) C.(–2, 0) D.(1, 2)

OpenStudy (mathstudent55):

You need to add the equations together to eliminate one variable. You may need to first multiply one or both equations by factors to eliminate a variable. Look at both equations, if you add them, you'll get 7x + y = -2. No variable was eliminated. That means you need to multiply. Let's eliminate the x variable. 1. Multiply the first equation by -5 and the second equation by 2: -5(2x + 6y) = -5(-12) 2(5x - 5y) = 2(10) -10x - 30y = 60 10x - 10y = 20 Now add them: -40y = 80 Solve: y = -2 You can now plug in y = -2 into one of the original equations to get x.

OpenStudy (shaik0124):

x=0 so b is answer

OpenStudy (mathstudent55):

Correct